OpenStudy (anonymous):

Factoring : 4y^2-8y-2xy

OpenStudy (cwrw238):

first find common factor of 3 terms (the GCF) GCF of 4,8 and 2 = 2 GCF of y^2, y and xy is y so out gcf is 2y can you continue?

OpenStudy (anonymous):

Yes. I will post my answer later on here, mind checking it then?

OpenStudy (cwrw238):

no problem

OpenStudy (anonymous):

I got 4y^2-8y-2xy!

OpenStudy (cwrw238):

thats the original expression you asked to factorize it : GCF = 2y now you need to divide each term by the GCF 4y^2 / 2y = 2y -8y / 2y = -4 -2xy / 2y = ? can you do the last one?

OpenStudy (anonymous):

No, I know. I copied the wrong one! LOL, my bad! So silly of me, but that is what I got. :)

OpenStudy (anonymous):

Mind helping me with some others?

OpenStudy (cwrw238):

you sure you can finish this one? need more help?

OpenStudy (anonymous):

I got 2y(2y-4-x)

OpenStudy (cwrw238):

yea

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

2ax+2bx-3a-3b, how would I start off on that one?

OpenStudy (cwrw238):

here there is no common factor of all the terms so try grouping: 2ax + 2bx - 3a - 3b = 2x(a + b) - 3(a + b) (a - b) is a common factor:, so its (2x - 3)(a + b)

OpenStudy (cwrw238):

* (a + b) is a common factor

OpenStudy (anonymous):

Okay, thanks. I am going to try to do a few of my own and then have you check them? :)

OpenStudy (cwrw238):

thats the way to go

OpenStudy (anonymous):

Can you help me with C^3-27?